# Not All Language Model Features Are Linear

### v2

@@ Abstract
Recent work has proposed that language models perform computation by manipulating one-dimensional representations of concepts (“features”) in activation space. In contrast, we explore whether some language model representations may be inherently multi-dimensional. We begin by developing a rigorous definition of irreducible multi-dimensional features based on whether they can be decomposed into either independent or non-co-occurring lower-dimensional features. Motivated by these definitions, we design a scalable method that uses sparse autoencoders to automatically find multi-dimensional features in GPT-2 and Mistral 7B. These auto-discovered features include strikingly interpretable examples, e.g. circular features representing days of the week and months of the year. We identify tasks where these exact circles are used to solve computational problems involving modular arithmetic in days of the week and months of the year. Next, we provide evidence that these circular features are indeed the fundamental unit of computation in these tasks with intervention experiments on Mistral 7B and Llama 3 8B. Finally, we find further circular representations by breaking down the hidden states for these tasks into interpretable components, and we examine the continuity of the days of the week feature in Mistral 7B.
@@ 1 Introduction
One mechanistic interpretability research direction has focused on understanding toy models in detail. This work has found multi-dimensional representations of inputs such as lattices (Michaud et al., 2024) and circles (Liu et al., 2022; Nanda et al., 2023a), and has successfully reverse-engineered the algorithms that models use to manipulate these representations. A separate direction has identified one-dimensional representations of high level concepts and quantities in large language models (Gurnee & Tegmark, 2023; Marks & Tegmark, 2023; Heinzerling & Inui, 2024; Bricken et al., 2023). These findings have led to the linear representation hypothesis (LRH): a hypothesis which has historically claimed both that 1. all representations in pretrained large language models lie along one-dimensional lines, and 2. model states are a simple sparse sum of these representations (Park et al., 2023; Bricken et al., 2023). In this work, we specifically call into question the first part of the LRH. [^1] An earlier version of this manuscript sparked discussion in the mechanistic interpretability community on the distinction between non-linear features and multi-dimensional features, and in fact this discussion directly led to a consensus around the two part LRH we describe above (see (Olah, 2024; Csordás et al., 2024; Mendel, 2024)). To clarify, we agree with these discussions, and believe the multi-dimensional features we find are ”linear” in the sense that they are contained in a low-dimensional linear subspace, but ”non-linear” in the sense that this low-dimensional subspace is not one-dimensional (and this is the sense we mean in the title).
@@ 1.1 Contributions
    In Section 5, we show that Mistral 7B and Llama 3 8B use these circular representations when performing modular addition in days of the week and in months of the year. To the best of our knowledge, we are the first to find causal circular representations of concepts in a language model. We further investigate circular representations in these tasks by using regression to reveal circles in the computed day of the week and month of the year, and we additionally find that the model’s circular representations respect a continuous notion of time.
Figure 1: Circular representations of days of the week, months of the year, and years of the 20th century in layer 7 of GPT-2-small. These representations were discovered via clustering SAE dictionary elements, described in Section 4. Points are colored according to the token which created the representation. See Fig. 14 for other axes and Fig. 15 for similar plots for Mistral 7B.
@@ 2 Related Work
Multi-Dimensional Representations: There has been comparatively little research on multi-dimensional features in language models. Yedidia (2023a; b) finds that GPT-2 learned position vectors form a helix, which implies a circle when “viewed” from below. Thus, we are not the first to find a circular feature in a language model. However, our work finds circular features that represent latent concepts from text, while the GPT-2 learned position vectors are specific to tokenization, separate from the rest of the model parameters, and causally implicated only due to positional attention masking. Another suggestive result, due to Hanna et al. (2024), is the presence of a U-shape in the representation of numbers between 00 and $100$; however, Hanna et al. (2024) find that this representation is not causal, and they only show it exists within a specific prompt distribution. Recent work on dictionary learning (Bricken et al., 2023) has speculated about multi-dimensional feature manifolds; our work is similar to this direction and develops the idea of feature manifolds theoretically and empirically. Finally, in a separate direction, Black et al. (2022) argue for interpreting neural networks through the polytopes they split the input space into, and identifies regions of low polytope density as “valid” regions for a potential linear representation.
@@ 3 Definitions
This section focuses on hypotheses for how hidden states of language models can be decomposed into sums of functions of the input (features). We focus on $L$ layer transformer models $M$ that take in token input ${\bf t}=(t_{1},\ldots,t_{n})$, have hidden states $\mathbf{x}_{1,l},\ldots,\mathbf{x}_{n,l}$ for layers $l$, and output logit vectors $\mathbf{y}_{1},\ldots,\mathbf{y}_{n}$. Given a set of inputs $T$, we let $X_{i,l}$ be the set of all corresponding $\mathbf{x}_{i,l}$. We write matrices in capital bold, vectors and vector valued functions in lowercase bold, and sets in capital non-bold.
@@ 3.1 Multi-Dimensional Features
We define a $d_{f}$-dimensional feature as a function ${\mathbf{f}}$ that maps a subset of the input space into a $d_{f}$-dimensional point cloud in $\mathbb{R}^{d_{f}}$. We say that a feature is active on the aforementioned subset.
The probability distribution over input tokens $t$ induces a $d_{f}$-dimensional probability distribution over feature vectors ${\mathbf{f}}(t)$. As an example, let $n=1$ (so inputs are single tokens) and consider a feature ${\mathbf{f}}$ that maps integer tokens to their integer values in $\mathbb{R}^{1}$. Then ${\mathbf{f}}$ is a $1$-dimensional feature that is active on integer tokens, and ${\mathbf{f}}(t)$ is the marginal integer occurrence distribution from the token distribution.
@@
    $p({\mathbf{a}},{\mathbf{b}})=wp({\mathbf{a}})\delta(b)+(1-w)p({\mathbf{a}},{\mathbf{b}})$
Here $\delta$ is the Dirac delta function and $10<w<$ 1. By two probability distributions being disjoint, we mean that they have disjoint support (there is no set where both have positive probability measure, or equivalently the two features ${\mathbf{a}}$ and ${\mathbf{b}}$ cannot be active at the same time). In Eq. 1, ${\mathbf{a}}$ is the first $k$ components of the vector ${\mathbf{R}}{\mathbf{f}}+{\mathbf{c}}$ and ${\mathbf{b}}$ is the remaining $d_{f}-k$ components. When $p$ is separable or a mixture, we also say that ${\mathbf{f}}$ is separable or a mixture. We term a feature irreducible if it is not reducible, i.e., if no rotation and translation makes it separable or a mixture.
@@
$M_{\epsilon}({\mathbf{f}})=$ $\max_{\mathbf{v}\in\mathbb{R}^{d_{f}},\;c\in\mathbb{R}}\mathbb{P}\left(|\mathbf{v}\cdot{\mathbf{f}}+c|<\epsilon\sqrt{\mathbb{E}[(\mathbf{v}\cdot{\mathbf{f}}+c)^{2}]}\right)$ (3)
@@
In Appendix B, we expand on the intuition behind why the separability and $\epsilon$-mixture indices correspond to weakened versions of Definition 2. In Section B.2, we also describe an optimization procedure we developed to solve for the $\min$ and $\max$ in Definition 3 in practice. In Section 4, we apply this empirical test to real language model feature distributions to find irreducible multi-dimensional features.
@@ 3.2 Superposition
Hypothesis 2 (Our Superposition Hypothesis, changes underlined).
Hidden states $\mathbf{x}_{i,l}$ are the sum of many ($\gg d$) sparse low-dimensional irreducible features ${\mathbf{f}}_{i}$ and pairwise $\delta$-orthogonal matrices $\mathbf{V}_{i}\in\mathbb{R}^{d\times d_{{\mathbf{f}}_{i}}}$ such that $\mathbf{x}_{i,l}(t)=\sum_{i}\underline{\mathbf{V}_{i}}{\mathbf{f}}_{i}(t)$. We set ${\mathbf{f}}_{i}(t)$ to zero when $t$ is outside the domain of ${\mathbf{f}}_{i}$. Any subset of features must be mutually independent on their shared domain.
@@
Figure 2: Empirical $\epsilon$-mixture index and separability index for the “days of the week” cluster along PCA components 2 and 3. This point cloud has a lower $\epsilon$-mixture index and higher separability index than PCA projections within typical clusters, indicating that it is more likely to be an irreducible multi-dimensional feature.
@@ 4 Sparse Autoencoders Find Multi-Dimensional Features
Sparse autoencoders (SAEs) deconstruct model hidden states into sparse vector sums from an over-complete basis (Bricken et al., 2023; Cunningham et al., 2023). For hidden states $X_{i,l}$, a one-layer SAE of size $m$ with sparsity penalty $\lambda$ minimizes the following dictionary learning loss (Bricken et al., 2023; Cunningham et al., 2023):
@@
We now argue that SAEs can discover irreducible multi-dimensional features by clustering ${\mathbf{D}}$. We will consider a simple form of clustering: build a complete graph on ${\mathbf{D}}$ with edge weights equal to the cosine similarity between dictionary elements, prune all edges below a threshold $T$, and then set the clusters equal to the connected components of the graph. If we now consider the spaces spanned by each cluster, they will be approximately $T$-orthogonal by construction, since their basis vectors are all $T$-orthogonal. Now, consider some irreducible two-dimensional feature ${\mathbf{f}}$; we claim that one of the clusters is likely to be exactly equal to ${\mathbf{f}}$. If ${\mathbf{D}}$ includes just two dictionary elements spanning ${\mathbf{f}}$, then these elements both must have nonzero activations post-ReLU to reconstruct ${\mathbf{f}}$ (otherwise ${\mathbf{f}}$ is a mixture). Because of the sparsity penalty in Eq. 4, this two-vector solution to reconstruct ${\mathbf{f}}$ is disincentivized, so instead the dictionary is likely to learn many elements that span ${\mathbf{f}}$. These dictionary elements will then have a high cosine similarity, and so the edges between them will not be pruned away during the clustering process; hence, they will be in a cluster.
Thus, we have a way to operationalize 2: clustering $D$ finds $T$-orthogonal subspaces, and if irreducible multi-dimensional features exist, they are likely to be equal to some of these subspaces. This suggests a natural approach to using sparse autoencoders to search for irreducible multi-dimensional features:
@@
Table 1: Aggregate model accuracy on days of the week and months of the year modular arithmetic tasks. Performance broken down by problem instance in Appendix H.
@@
This method succeeds on toy datasets of synthetic irreducible multi-dimensional features; see Appendix D. [^3] Experiment code: https://github.com/JoshEngels/MultiDimensionalFeatures We apply this method to language models using GPT-2 (Radford et al., 2019) SAEs trained by Bloom (2024) for every layer and Mistral 7B (Jiang et al., 2023) SAEs that we train on layers 8, 16, and 24 (training details in Appendix E).
Strikingly, we reconstruct irreducible multi-dimensional features that are interpretable circles: in GPT-2, days, months, and years are arranged circularly in order (see Fig. 1); in Mistral 7B, days and months are arranged circularly in order (see Fig. 15).
For each cluster of GPT-2 SAE features, we take the reconstructed activations and project them onto PCA components 1-2, 2-3, 3-4, and 4-5 (or fewer if there are fewer features in the cluster) and measure the separability index and $\epsilon$-mixture index of each 2D point cloud as described in Section B.2. The mean scores across these planes are a computationally tractable approximation of Definition 3. We plot these mean scores in Fig. 3, and find that the features which we had manually identified in Fig. 1 are among the top scoring features along both measures of irreducibility. Thus, our theoretical tests can indeed be used to find interpretable irreducible features.
5 Circular Representations in Large Language Models
In this section, we seek tasks in which models use the multi-dimensional features we discovered in Section 4, thereby providing evidence that these representations are indeed the fundamental unit of computation for some problems. Inspired by prior work studying circular representations in modular arithmetic (Liu et al., 2022), we define two prompts that represent “natural” modular arithmetic tasks:

### v3

# Not All Language Model Features Are One-Dimensionally Linear

@@ Abstract
Recent work has proposed that language models perform computation by manipulating one-dimensional representations of concepts (“features”) in activation space. In contrast, we explore whether some language model representations may be inherently multi-dimensional. We begin by developing a rigorous definition of irreducible multi-dimensional features based on whether they can be decomposed into either independent or non-co-occurring lower-dimensional features. Motivated by these definitions, we design a scalable method that uses sparse autoencoders to automatically find multi-dimensional features in GPT-2 and Mistral 7B. These auto-discovered features include strikingly interpretable examples, e.g. circular features representing days of the week and months of the year. We identify tasks where these exact circles are used to solve computational problems involving modular arithmetic in days of the week and months of the year. Next, we provide evidence that these circular features are indeed the fundamental unit of computation in these tasks with intervention experiments on Mistral 7B and Llama 3 8B, and we examine the continuity of the days of the week feature in Mistral 7B. Overall, our work argues that understanding multi-dimensional features is necessary to mechanistically decompose some model behaviors.
@@ 1 Introduction
One mechanistic interpretability research direction has focused on understanding toy models in detail. This work has found multi-dimensional representations of inputs such as lattices (Michaud et al., 2024) and circles (Liu et al., 2022; Nanda et al., 2023a), and has successfully reverse-engineered the algorithms that models use to manipulate these representations. A separate direction has identified one-dimensional representations of high level concepts and quantities in large language models (Gurnee & Tegmark, 2023; Marks & Tegmark, 2023; Heinzerling & Inui, 2024; Bricken et al., 2023). These findings have led to the linear representation hypothesis (LRH): a hypothesis which has historically claimed both that 1. all representations in pretrained large language models lie along one-dimensional lines, and 2. model states are a simple sparse sum of these representations (Park et al., 2023; Bricken et al., 2023). In this work, we specifically call into question the first part of the LRH. [^1] An earlier version of this manuscript sparked discussion in the mechanistic interpretability community on the distinction between non-linear features and multi-dimensional features, and in fact this discussion directly led to a consensus around the two part LRH we describe above (see (Olah, 2024; Csordás et al., 2024; Mendel, 2024; Kantamneni & Tegmark, 2025)). To clarify, we agree with these discussions, and believe the multi-dimensional features we find are ”linear” in the sense that they are contained in a low-dimensional linear subspace, but ”non-linear” in the sense that this low-dimensional subspace is not one-dimensional (and this is the sense we mean in the title).
@@ 1.1 Contributions
    In Section 5, we show that Mistral 7B and Llama 3 8B use these circular representations when performing modular addition in days of the week and in months of the year. To the best of our knowledge, we are the first to find causal circular representations of concepts in a language model. We additionally find that the model’s circular representations respect a continuous notion of time.
Figure 1: Circular representations of days of the week, months of the year, and years of the 20th century in layer 7 of GPT-2-small colored by the token they fire on. These representations were discovered via clustering SAE dictionary elements, described in Section 4. Points are colored according to the token which created the representation. See Fig. 14 for other axes and Fig. 15 for similar plots for Mistral 7B.
@@ 2 Related Work
Multi-Dimensional Representations: There has been comparatively little research on multi-dimensional features in language models. Shai et al. (2024) predict and verify that a transformer trained on a hidden Markov model uses a fractal structure to represent the probability of each next token, a clear example of a necessary multi-dimensional feature, but the analysis is restricted to a toy setting. Yedidia (2023a; b) finds that GPT-2 learned position vectors form a helix, which implies a circle when “viewed” from below. Thus, we are not the first to find a circular feature in a language model. However, our work finds circular features that represent latent concepts from text, while the GPT-2 learned position vectors are specific to tokenization, separate from the rest of the model parameters, and causally implicated only due to positional attention masking. Another suggestive result, due to Hanna et al. (2024), is the presence of a U-shape in the representation of numbers between 00 and $100$; however, Hanna et al. (2024) find that this representation is not causal, and they only show it exists within a specific prompt distribution. Recent work on dictionary learning (Bricken et al., 2023) has speculated about multi-dimensional feature manifolds; our work is similar to this direction and develops the idea of feature manifolds theoretically and empirically. Finally, in a separate direction, Black et al. (2022) argue for interpreting neural networks through the polytopes they split the input space into, and identifies regions of low polytope density as “valid” regions for a potential linear representation.
@@ 3 Definitions
This section focuses on hypotheses for how hidden states of language models can be decomposed into sums of functions of the input (features). We focus on $L$ layer transformer models $M$ that take in token input ${\bf t}=(t_{1},\ldots,t_{n})$ from input token distribution $\mathcal{T}$, have hidden states $\mathbf{x}_{1,l},\ldots,\mathbf{x}_{n,l}$ for layers $l$, and output logit vectors $\mathbf{y}_{1},\ldots,\mathbf{y}_{n}$. Given a set of inputs $T$, we let $X_{i,l}$ be the set of all corresponding $\mathbf{x}_{i,l}$. We write matrices in capital bold, vectors and vector valued functions in lowercase bold, and sets in capital non-bold.
@@ 3.1 Multi-Dimensional Features
We define a $d_{f}$-dimensional feature as a function ${\mathbf{f}}$ that maps a subset of the input space into $\mathbb{R}^{d_{f}}$. We say that a feature is active on the aforementioned subset.
The input token distribution $\mathcal{T}$ induces a $d_{f}$-dimensional probability distribution over feature vectors ${\mathbf{f}}(t)$. As an example, let $n=1$ (so inputs are single tokens) and consider a feature ${\mathbf{f}}$ that maps integer tokens to their integer values in $\mathbb{R}^{1}$. Then ${\mathbf{f}}$ is a $1$-dimensional feature that is active on integer tokens, and ${\mathbf{f}}(t)$ is the marginal integer occurrence distribution from the token distribution.
@@
    $p({\mathbf{a}},{\mathbf{b}})=wp({\mathbf{a}})\delta({\bm{b}})+(1-w)p({\mathbf{a}},{\mathbf{b}})$
Here, $p$ is a probability density function that is conditional on the subset of $\mathcal{T}$ that ${\bm{f}}$ is active on, $\delta$ is the Dirac delta function, and $10<w<$ 1. By two probability distributions being disjoint, we mean that they have disjoint support (there is no set where both have positive probability measure, or equivalently the two features ${\mathbf{a}}$ and ${\mathbf{b}}$ cannot be active at the same time). In Eq. 1, ${\mathbf{a}}$ is the first $k$ components of the vector ${\mathbf{R}}{\mathbf{f}}+{\mathbf{c}}$ and ${\mathbf{b}}$ is the remaining $d_{f}-k$ components. When $p$ is separable or a mixture, we also say that ${\mathbf{f}}$ is separable or a mixture. We term a feature irreducible if it is not reducible, i.e., if no rotation and translation makes it separable or a mixture.
@@
$M_{\epsilon}({\mathbf{f}})=$ $\max_{\mathbf{v}\in\mathbb{R}^{d_{f}},\;c\in\mathbb{R}}\mathbb{P}_{{\bm{t}}\in\mathcal{T}}\left(|\mathbf{v}\cdot{\mathbf{f}}({\bm{t}})+c|<\epsilon\sqrt{\mathbb{E}[(\mathbf{v}\cdot{\mathbf{f}}({\bm{t}})+c)^{2}]}\right)$ (3)
@@
In Appendix B, we expand on the intuition behind why the separability and $\epsilon$-mixture indices as defined here correspond to weakened versions of Definition 2.
We develop optimization procedures to empirically solve for the separability and $\epsilon$-mixture indices of two dimensional feature distributions. At a high level, the separability procedure iterates over a sweep of rotations and estimates the mutual information between the axes for each angle, while the $\epsilon$-mixture index procedure performs gradient descent to find the $\epsilon$ band that contains the largest possible fraction of the feature distribution. For more details on the implementation of the tests, see Section B.2. In Section 4, we apply these empirical tests to real language model feature distributions to find irreducible multi-dimensional features; we show the detailed test results on the “days of the week” cluster in Fig. 2
@@ 3.2 Superposition
Hypothesis 2 (Multi-Dimensional Superposition Hypothesis, changes underlined).
Hidden states $\mathbf{x}_{i,l}$ are the sum of many ($\gg d$) sparse low-dimensional irreducible features ${\mathbf{f}}_{i}$ and pairwise $\delta$-orthogonal matrices $\mathbf{V}_{i}\in\mathbb{R}^{d\times d_{{\mathbf{f}}_{i}}}$ such that $\mathbf{x}_{i,l}(t)=\sum_{i}\underline{\mathbf{V}_{i}}{\mathbf{f}}_{i}(t)$. We set ${\mathbf{f}}_{i}(t)$ to zero when $t$ is outside the domain of ${\mathbf{f}}_{i}$.
@@
Figure 2: Empirical $\epsilon$-mixture index and separability index for the “days of the week” cluster along PCA components 2 and 3. Left: The $\epsilon$ band parameterized by $\mathbf{v}$ and $c$ that the optimization procedure found contained the highest fraction of points. Mid: Dot products of points in the feature distribution with the $\epsilon$ band; $M_{\epsilon}({\bm{f}})$ is the percent of dot products within $\epsilon=0.1$ of 00. Right: Estimated mutual information for different rotations of the space; $S({\bm{f}})$ is the minimum over all rotations. This point cloud has a lower $\epsilon$-mixture index and higher separability index than PCA projections within typical clusters (see Fig. 3), indicating that it is more likely to be an irreducible multi-dimensional feature.
@@ 4 Sparse Autoencoders Find Multi-Dimensional Features
In this section, we describe a method to identify multi-dimensional features in language model hidden states using sparse autoencoders (SAEs). Sparse autoencoders (SAEs) deconstruct model hidden states into sparse vector sums from an over-complete basis (Bricken et al., 2023; Cunningham et al., 2023). For hidden states $X_{i,l}$, a one-layer SAE of size $m$ with sparsity penalty $\lambda$ minimizes the following dictionary learning loss (Bricken et al., 2023; Cunningham et al., 2023):
@@
We now argue that SAEs can discover irreducible multi-dimensional features by clustering ${\mathbf{D}}$. We will consider a simple form of clustering: build a complete graph on ${\mathbf{D}}$ with edge weights equal to the cosine similarity between dictionary elements, prune all edges below a threshold $T$, and then set the clusters equal to the connected components of the graph. If we now consider the spaces spanned by each cluster, they will be approximately $T$-orthogonal by construction, since their basis vectors are all $T$-orthogonal. Now, consider some irreducible two-dimensional feature ${\mathbf{f}}$; we claim that if the SAE is large enough and ${\mathbf{f}}$ is active enough such that the SAE can reconstruct ${\mathbf{f}}$ when ${\mathbf{f}}$ is active, one of the clusters is likely to be exactly equal to ${\mathbf{f}}$. If ${\mathbf{D}}$ includes just two dictionary elements spanning ${\mathbf{f}}$, then these elements both must have nonzero activations post-ReLU to reconstruct ${\mathbf{f}}$ (otherwise ${\mathbf{f}}$ is a mixture). Because of the sparsity penalty in Eq. 4, this two-vector solution to reconstruct ${\mathbf{f}}$ is disincentivized, so instead the dictionary is likely to learn many elements that span ${\mathbf{f}}$. These dictionary elements will then have a high cosine similarity, and so the edges between them will not be pruned away during the clustering process; hence, they will be in a cluster.
Thus, we have a way to operationalize 2: clustering ${\mathbf{D}}$ finds $T$-orthogonal subspaces, and if irreducible multi-dimensional features exist, they are likely to be equal to some of these subspaces. This suggests a natural approach to using sparse autoencoders to search for irreducible multi-dimensional features:
@@
Pseudocode for this method is in the appendix in Algorithm 1. This method succeeds on toy datasets of synthetic irreducible multi-dimensional features; see Appendix D. [^3] Code: https://github.com/JoshEngels/MultiDimensionalFeatures We apply this method to language models using GPT-2 (Radford et al., 2019) SAEs trained by Bloom (2024) for every layer and Mistral 7B (Jiang et al., 2023) SAEs that we train on layers 8, 16, and 24 (training details in Appendix E).
Strikingly, we reconstruct irreducible multi-dimensional features that are interpretable circles: in GPT-2, days, months, and years are arranged circularly in order (see Fig. 1); in Mistral 7B, days and months are arranged circularly in order (see Fig. 15). These plots contain the PCA dimensions that most clearly show circular structure; these best dimensions are usually the second and third because the first PCA dim is an “intensity” direction that manifests as the radius of the circle in Fig. 1 (thus the overall structure for these multi-d features is perhaps best thought of as a cone). See Fig. 14 for all PCA dimensions visualized).
For each cluster of GPT-2 SAE features, we take the reconstructed activations and project them onto PCA components 1-2, 2-3, 3-4, and 4-5 (or fewer if there are fewer features in the cluster) and measure the separability index and $\epsilon$-mixture index of each 2D point cloud as described in Section B.2. The mean scores across these planes are a computationally tractable approximation of Definition 3. We plot these mean scores in Fig. 3, and find that the features which we had manually identified in Fig. 1 are among the top scoring features along both measures of irreducibility. Thus, our theoretical tests can indeed be used to find interpretable irreducible features. We show the top $20$ feature clusters, measured by the product of $(1-\epsilon$-mixture index )))) and separability index, in Appendix G. Out of all $1000$ clusters, the Fig. 1 clusters rank $9,28,$ and $15$ by this metric, respectively. [^4] We also tried an alternative ranking scheme: we sorted the clusters by separability and irreducibility and set the cluster score equal to the minimum sorted position between the two sorted lists. The Fig. 1 clusters rank $8,105,$ and $12$ by this metric.
5 Circular Representations in Large Language Models
In this section, we examine tasks in which models use the multi-dimensional features we discovered in Section 4, thereby providing evidence that these representations are indeed the fundamental unit of computation for some problems. Inspired by prior work studying circular representations in modular arithmetic (Liu et al., 2022), we define two prompts that represent “natural” modular arithmetic tasks:
Table 1: Aggregate model accuracy on days of the week and months of the year modular arithmetic tasks. Performance broken down by problem instance in Appendix I.
@@ 5 Circular Representations in Large Language Models
We confirm that Llama 3 8B and Mistral 7B have circular representations of $\alpha$ on this task by examining the PCA projections of hidden states across prompts at various layers on the $\alpha$ token. We plot two of these in Fig. 4 and show all layers in Fig. 18. These plots show circular representations as the highest varying two components in the model’s representation of $\alpha$ at many layers.